# Sensitivity of CTAO to axion-like particles from blazars: a machine learning approach

### v1

@@ 2.2 Modelling of ALP-photon oscillations
The propagating ALP-photon beam can be described as a three-state quantum system, representing the two photon polarizations and the ALP state, obeying a Schrödinger-like equation [raffelt1988]. However, in the general case of an unpolarized beam (as for gamma rays), the system is more correctly represented by a $3\times 3$ density matrix $\rho$ satisfying the von Neumann-like equation [deangelis2011, galanti2022]
@@
where $y$ is the propagation direction and $\mathcal{M}_{0}(E)$ is an energy-dependent matrix whose elements account for photon dispersion, absorption and for ALP-photon mixing in external magnetic fields. For a given magnetic field configuration, this equation can be solved iteratively over discrete magnetic domains by the method of transfer matrices [mirizzi2008, bassan2010, deangelis2011, galanti2022]. Given the initial state $\rho_{i}$ at position $y_{0}$ and the total transfer matrix $\mathcal{U}(E;y,y_{0})$, the probability of finding the system in the final state $\rho_{f}$ at position $y$ is expressed as
@@
In this work, the calculations outlined above are performed using the gammaALPs v0.3.0333https://github.com/me-manu/gammaALPs Python package [meyer2021]. This package defines a ModuleList class where different magnetic field environments can be added in sequence. The ModuleList can then be run to automatically solve eq. (2.4) over the line of sight from a given source and obtain $P_{\gamma\gamma}(E)$ following eq. (2.6), for initial state $\rho_{0}$ and fixed ALP parameters $(m_{a},g_{a\gamma})$.
In order to compute the photon survival probabilities for our selected blazars, we consider ALP-photon mixing in their jets’ magnetic fields, modelled with helical and tangled components as in ref. [davies2021]. The model is based on the blazar jet structure described by Potter and Cotter [potter2013a, potter2013b, potter2013c, potter2013d] and implemented in the gammaALPs class MixJetHelicalTangled. The parameters of each source are adapted from Figure 7 in ref. [potter2015] [^4] We note that ref. [potter2015] employs the same source nomenclature used in the the Fermi-LAT Bright Source List (0FGL) [abdo2009, abdo2010], and hence the source PKS 2155-304 appears listed as J2158., except for the free electron density normalization which is taken from Table 5 in ref. [tavecchio2010]. Additionally, we include photon-ALP mixing in the Milky Way’s galactic magnetic field as modelled in ref. [jansson2012] and EBL absorption from ref. [dominguez2011]. These propagation environments are implemented, respectively, in the classes MixGMF and OptDepth (from the ebltable package [^5] https://github.com/me-manu/ebltable). Finally, a possible contribution from the intergalactic magnetic field is usually modeled as a randomly oriented field with present-day intensity $B_{0}\lesssim 1\,{\rm nG}$ over a coherence length $l_{0}\lesssim 1\,{\rm Mpc}$ [ade2016]. While the effect of this field is not necessarily small (e.g [deangelis2007, abramowski2013]), we decided to neglect it owing to the large uncertainty in its structure and magnitude, in line with previous works [meyer2014b, ajello2016, abdalla2021, abe2024].
@@
Figure 1: ALP effects on the baseline and flaring spectral shapes chosen for the sources in this work. For large enough values of $g_{a\gamma}$, both spectral distortions (“wiggles”) and a hardening at high energies are visible.
@@ 2.3 Data simulation and analysis
CTAO observations are simulated using the Gammapy v1.0.1 Python package [gammapy2023, gammapy-1.0.1]. For each source, we generate 50 hours of mock data for the baseline state and 5 hours (corresponding to about a night of observations) for the flaring state. We assume observations to be taken in the ON-OFF mode, i.e. by pointing the telescopes at an offset from the source (the ON region) and estimating the background from a number of OFF regions symmetrically distributed around the center of the field of view. For this setup, we consider an offset from the source of 0.7 degrees, taking ON and OFF regions with a radius of 0.2 degrees. We choose 3 OFF regions, or equivalently, an acceptance ratio $\alpha=1/3$. The source coordinates are implemented in the SkyCoord and SourceCatalog4FGL classes, respectively from the Astropy [astropy2013, astropy2018, astropy2022] and Gammapy [^6] Note that Gammapy v1.3 [gammapy-1.3] is necessary to load the 4FGL-DR4 catalog. packages. The parameters of flaring spectra are taken directly from refs. [albert2007, aharonian2009], and the redshift parameters reported in TeVCat [^7] https://www.tevcat.org/ [wakely2008] are used.
We subdivide the measured energy axis with 20 bins per energy decade, in order to resolve the fine spectral irregularities that may be introduced by ALP effects. For each energy bin $i$, the expected number of signal counts $\mu_{i}$ is obtained by convolving the observed flux $\phi_{\rm obs}$ with the instrument response functions (IRFs), and then integrating over the energy bin width and observation time [abdalla2021]. For our simulations we use the publicly available CTAO prod5 v0.1 IRFs [ctao_irfs_prod5], computed for the CTAO “Alpha” configuration described in Section 1. These include effective area, point spread function (PSF) and energy dispersion, as well as the background rates that can be used to compute the expected number of background counts per energy bin $b_{i}$. Different IRF files are available for 0.5, 5 and 50 hours of observations, both for the Northern and Southern Arrays in distinct subarray configurations, and for different values of zenith and azimuth angles. We simulate observations of Mrk 501 and PKS 2155-304 with the Northern and Southern Arrays, respectively, assuming all observations to be taken at a zenith angle of 20 degrees while averaging over the azimuth angle. The astrophysical and observational parameters used in simulating the datasets, as well as the chosen IRFs and spectral models, are summarised in Tables 1 and 2.
@@
Table 1: Summary table of the astrophysical parameters of the selected blazars. The gammaALPs parameters are adapted from refs. [potter2015] and [tavecchio2010]. Additionally, the parameters rvhe and r_T are used, both set equal to r_0 and taken to be $\sim 100$ times the size of the transition region $R_{T}$. Unspecified parameters are set to their default values.
Source Livetime [h] IRFs Reference spectrum Mrk 501 (baseline) 50 Prod5-North-20deg-AverageAz-4LSTs09MSTs.180000s LP [4fgl-dr4] PKS 2155-304 (baseline) 50 Prod5-South-20deg-AverageAz-14MSTs37SSTs.180000s LP [4fgl-dr4] Mrk 501 (flaring) 5 Prod5-North-20deg-AverageAz-4LSTs09MSTs.18000s LP [albert2007] PKS 2155-304 (flaring) 5 Prod5-South-20deg-AverageAz-14MSTs37SSTs.18000s ECPL [aharonian2009]
Table 2: Observation times, IRF file names and reference spectra for the selected blazars.
@@
Figure 2: Example simulated SEDs for two 50-hour observations of the Mrk 501 baseline state. Spectral fits are shown for the ALP-less case and for two different ALP parameter sets. The residuals for both cases are computed as $(f-\phi)/\sigma_{f}$, where $f$ is the measured flux with uncertainty $\sigma_{f}$ and $\phi(E)$ is the model prediction (see ref. [abdalla2021]).
@@ 3 Derivation of ALP sensitivity with the likelihood-ratio test
We consider an array of $10\times 10$ logarithmically spaced points over the ALP parameter space, with $m_{a}\in[0.1,1000]\,{\rm neV}$ and $g_{a\gamma}\in[0.03,7]\times 10^{-11}\,{\rm GeV}^{-1}$. The choice of the mass interval is commonly adopted in similar studies, and roughly corresponds to the region where the critical energy for the strong mixing regime falls within the CTAO energy range [meyer2014a]. Such energy value is defined as that above which the ALP-photon oscillation probability in a constant magnetic field becomes maximal and almost independent of energy [meyer2014a, mirizzi2008, bassan2010, deangelis2011, galanti2022]. On the other hand, the upper bound on $g_{a\gamma}$ is motivated by the most recent experimental limit set by the CAST helioscope [altenmuller2024].
Given the intrinsic spectrum of each source, we compute the photon survival probability $P_{\gamma\gamma}$ for each $(m_{a},g_{a\gamma})$ pair with gammaALPs as described in Section 2.2, and obtain the expected observed spectrum as in eq. (2.7).
$\mathcal{L}(m_{a},g_{a\gamma},\mu,b,\theta|D)=\prod_{i}\mathcal{L}(m_{a},g_{a\gamma},\mu_{i},b_{i},\theta|D_{i})$ (3.1)
@@
$\mathcal{L}(m_{a},g_{a\gamma},\mu_{i},b_{i},\theta|D)=\frac{(\mu_{i}+\alpha\,b_{i})^{n_{{\rm ON},i}}}{n_{{\rm ON},i}!}e^{-(\mu_{i}+\alpha\,b_{i})}\frac{b_{i}^{n_{{\rm OFF},i}}}{n_{{\rm OFF},i}!}e^{-b_{i}}$ (3.2)
@@
Following ref. [abe2024], we then define the test statistic
${\rm TS}(m_{a},g_{a\gamma}|D)=-2\ln\frac{\mathcal{L}(m_{a},g_{a\gamma},\hat{\mu},\hat{b},\hat{\theta}|D)}{\hat{\mathcal{L}}(D)}\,$, (3.3)
@@
${\rm WSTAT}=-2\ln\mathcal{L}=2\sum_{i}[\mu_{i}+(\alpha+1)b_{i}-n_{{\rm ON},i}\ln(\mu_{i}+\alpha b_{i})-n_{{\rm OFF},i}\ln b_{i}]\,$, (3.4)
where terms that do not depend on $\mu_{i}$ or $b_{i}$ have been neglected. This quantity is computed and returned by Gammapy when fitting each spectral model to the dataset $D$. Therefore, eq. (3.3) is equivalent to
${\rm TS}(m_{a},g_{a\gamma}|D)={\rm WSTAT}(m_{a},g_{a\gamma}|D)-{\rm WSTAT}_{\rm min}(D)\,.$ (3.5)
@@
Figure 3: Example likelihood-ratio TS distribution obtained for the baseline state of Mrk 501.
For each source (considered in both baseline and flaring states) we simulate 100 observations without including ALP effects, following the prescriptions outlined in Section 2.3, and compute the average TS for each $(m_{a},g_{a\gamma})$ pair. In order to obtain the statistical significance for excluding the ALP hypothesis with parameters $(m_{a},g_{a\gamma})$, one then needs to compare the obtained average, denoted with ${\rm TS_{0}}(m_{a},g_{a\gamma})$, against the TS distribution expected under the ALP hypothesis. As has been argued several times, Wilks’ theorem [wilks1938] cannot be applied to the case of ALPs [meyer2014b, abe2024, abdalla2021], and thus the shape of the TS distribution cannot be assumed a priori. Therefore, for each $(m_{a},g_{a\gamma})$ point and each considered source, we simulate 100 datasets including ALP effects and compute the corresponding TS values according to eq. (3.5). An example of a TS distribution obtained in this way is shown in Figure 3. Following ref. [abe2024], we fit this distribution with a Gamma probability distribution function (PDF), given by
$f(x,\alpha,\beta)=\frac{\beta^{\alpha}x^{\alpha-1}e^{-\beta x}}{\Gamma(\alpha)}\,$, (3.6)
@@
All the statistical computations discussed above were performed using the Scipy Python package [virtanen2020].
@@ 4.1 Machine learning classifiers
A basic example of a ML classifier is the binary decision tree (BDT). A BDT can act on a dataset [^8] In machine learning literature the term “sample” would be used, whereas “dataset” would denote a set of many samples., identified by a vector of features $\mathbf{X}=(X_{1},\dots,X_{n})$, and distinguish whether it belongs to one of two classes by operating sequential binary splits on the features’ values. The splitting thresholds at each step (node) are referred to as the parameters of the tree; on the other hand, quantities such as the tree depth, maximum number of splits, etc. are called hyperparameters. During the “supervised training” of the BDT, the splits are chosen in order to minimize some loss function, by operating on datasets whose class is known a priori. Once the tree has been trained and its parameters have been fixed, it can be used to classify previously “unseen” test datasets. There are several methods to evaluate the performance of a ML classifier during the training phase, e.g. evaluating its accuracy and related metrics with train-test splitting and cross-validation. This ensures that a similar performance will be achieved on new data as well. A vast literature of textbooks and papers exists on the many scientific applications of ML classifiers and BDTs; see e.g. ref. [raschka2017] for a basic introduction.
Random forests (RFs) are an example of ensemble classifiers, since, as the name suggests, they are based on averaging a large number of binary decision trees. In order to avoid overfitting, i.e. classifiers performing “too well” on training data but badly on test data, it is necessary that the single BDTs stop after few splits (it is said that they are “weak” classifiers). BDTs, RFs and several other ML classification algorithms are implemented in the open-source Python package Scikit-learn [pedregosa2012]. When one of these algorithms is trained to distinguish between two classes and then applied to some sample, it typically returns a set of scores indicating how likely the given sample is to belong to each of the classes.
Machine learning classifiers such as RFs are widely used in Cherenkov telescope data analysis, specifically in steps that involve the so-called gamma-hadron separation [bock2004, albert2008, ohm2009]. Here, the atmospheric shower images are parametrized by a set of quantities (the Hillas parameters [hillas1985]) which are then used as features to distinguish between showers initiated by hadrons and by gamma-ray photons.
XGBoost (eXtreme Gradient Boosting) [chen2016] is a similar algorithm to the RF, which leverages a technique called gradient boosting to efficiently minimise the loss function. Implementations of the XGBoost algorithm are available in several programming languages, including Python [^9] https://xgboost.readthedocs.io/en/stable/, and generally exhibit reasonably good performance “out of the box”, with minimal tweaking compared to RFs.

### v2

@@ 2.2 Modelling of ALP-photon oscillations
The propagating ALP-photon beam can be described as a three-state quantum system, representing the two photon polarizations and the ALP state, obeying a Schrödinger-like equation Raffelt and Stodolsky (1988). However, in the general case of an unpolarized beam (as for gamma rays), the system is more correctly represented by a $3\times 3$ density matrix $\rho$ satisfying the von Neumann-like equation De Angelis et al. (2011); Galanti and Roncadelli (2022)
@@
where $y$ is the propagation direction and $\mathcal{M}_{0}(E)$ is an energy-dependent matrix whose elements account for photon dispersion, absorption and for ALP-photon mixing in external magnetic fields. For a given magnetic field configuration, this equation can be solved iteratively over discrete magnetic domains by the method of transfer matrices Mirizzi et al. (2008); Bassan et al. (2010); De Angelis et al. (2011); Galanti and Roncadelli (2022). Given the initial state $\rho_{i}$ at position $y_{0}$ and the total transfer matrix $\mathcal{U}(E;y,y_{0})$, the probability of finding the system in the final state $\rho_{f}$ at position $y$ is expressed as
@@
In this work, the calculations outlined above are performed using the gammaALPs v0.3.0333https://github.com/me-manu/gammaALPs Python package Meyer et al. (2021). This package defines a ModuleList class where different magnetic field environments can be added in sequence. The ModuleList can then be run to automatically solve eq. (5) over the line of sight from a given source and obtain $P_{\gamma\gamma}(E)$ following eq. (7), for initial state $\rho_{0}$ and fixed ALP parameters $(m_{a},g_{a\gamma})$.
In order to compute the photon survival probabilities for our selected blazars, we consider ALP-photon mixing in their jets’ magnetic fields, modelled with helical and tangled components as in ref. Davies et al. (2021). The model is based on the blazar jet structure described by Potter and Cotter Potter and Cotter (2012a, b, 2013a, 2013b) and implemented in the gammaALPs class MixJetHelicalTangled. The parameters of each source are adapted from Figure 7 in ref. Potter and Cotter (2015) [^4] We note that ref. Potter and Cotter (2015) employs the same source nomenclature used in the the Fermi-LAT Bright Source List (0FGL) Abdo et al. (2009, 2010), and hence the source PKS 2155-304 appears listed as J2158., except for the free electron density normalization which is taken from Table 5 in ref. Tavecchio et al. (2010). Additionally, we include photon-ALP mixing in the Milky Way’s galactic magnetic field as modelled in ref. Jansson and Farrar (2012) and EBL absorption from ref. Domínguez et al. (2011). These propagation environments are implemented, respectively, in the classes MixGMF and OptDepth (from the ebltable package [^5] https://github.com/me-manu/ebltable). Finally, a possible contribution from the intergalactic magnetic field is usually modeled as a randomly oriented field with present-day intensity $B_{0}\lesssim 1\,{\rm nG}$ over a coherence length $l_{0}\lesssim 1\,{\rm Mpc}$ Ade et al. (2016). While the effect of this field is not necessarily small (e.g De Angelis et al. (2007); Abramowski et al. (2013)), we decided to neglect it owing to the large uncertainty in its structure and magnitude, in line with previous works Meyer and Conrad (2014); Ajello et al. (2016); Abdalla et al. (2021); Abe et al. (2024a).
@@
Figure 1: ALP effects on the baseline and flaring spectral shapes chosen for the sources in this work. For large enough values of $g_{a\gamma}$, both spectral distortions (“wiggles”) and a hardening at high energies are visible
@@ 2.3 Data simulation and analysis
CTAO observations are simulated using the Gammapy v1.0.1 Python package Donath et al. (2023); Acero et al. (2023). For each source, we generate 50 hours of mock data for the baseline state and 5 hours (corresponding to about a night of observations) for the flaring state. We assume observations to be taken in the ON-OFF mode, i.e. by pointing the telescopes at an offset from the source (the ON region) and estimating the background from a number of OFF regions symmetrically distributed around the center of the field of view. For this setup, we consider an offset from the source of 0.7 degrees, taking ON and OFF regions with a radius of 0.2 degrees. We choose 3 OFF regions, or equivalently, an acceptance ratio $\alpha=1/3$. The source coordinates are implemented in the SkyCoord and SourceCatalog4FGL classes, respectively from the Astropy Astropy Collaboration and others (2013, 2018, 2022) and Gammapy [^6] Note that Gammapy v1.3 Acero et al. (2025) is necessary to load the 4FGL-DR4 catalog. packages. The parameters of flaring spectra are taken directly from refs. Albert et al. (2007); Aharonian et al. (2009), and the redshift parameters reported in TeVCat [^7] https://www.tevcat.org/ Wakely and Horan (2008) are used.
Table 1: Summary table of the astrophysical parameters of the selected blazars. The gammaALPs parameters are adapted from refs. Potter and Cotter (2015) and Tavecchio et al. (2010). Additionally, the parameters rvhe and r_T are used, both set equal to r_0 and taken to be $\sim 100$ times the size of the transition region $R_{T}$. Unspecified parameters are set to their default values
@@
Table 2: Observation times, IRF file names and reference spectra for the selected blazars
Source Livetime [h] IRFs Reference spectrum Mrk 501 (baseline) 50 Prod5-North-20deg-AverageAz-4LSTs09MSTs.180000s LP Ballet et al. (2024) PKS 2155-304 (baseline) 50 Prod5-South-20deg-AverageAz-14MSTs37SSTs.180000s LP Ballet et al. (2024) Mrk 501 (flaring) 5 Prod5-North-20deg-AverageAz-4LSTs09MSTs.18000s LP Albert et al. (2007) PKS 2155-304 (flaring) 5 Prod5-South-20deg-AverageAz-14MSTs37SSTs.18000s ECPL Aharonian et al. (2009)
Figure 2: Example simulated SEDs for two 50-hour observations of the Mrk 501 baseline state. Spectral fits are shown for the ALP-less case and for two different ALP parameter sets. The residuals for both cases are computed as $(f-\phi)/\sigma_{f}$, where $f$ is the measured flux with uncertainty $\sigma_{f}$ and $\phi(E)$ is the model prediction (see ref. Abdalla et al. (2021))
We subdivide the measured energy axis with 20 bins per energy decade, in order to resolve the fine spectral irregularities that may be introduced by ALP effects. For each energy bin $i$, the expected number of signal counts $\mu_{i}$ is obtained by convolving the observed flux $\phi_{\rm obs}$ with the instrument response functions (IRFs), and then integrating over the energy bin width and observation time Abdalla et al. (2021). For our simulations we use the publicly available CTAO prod5 v0.1 IRFs Cherenkov Telescope Array Observatory and Cherenkov Telescope Array Consortium (2021), computed for the CTAO “Alpha” configuration described in Section 1. These include effective area, point spread function (PSF) and energy dispersion, as well as the background rates that can be used to compute the expected number of background counts per energy bin $b_{i}$. Different IRF files are available for 0.5, 5 and 50 hours of observations, both for the Northern and Southern Arrays in distinct subarray configurations, and for different values of zenith and azimuth angles. We simulate observations of Mrk 501 and PKS 2155-304 with the Northern and Southern Arrays, respectively, assuming all observations to be taken at a zenith angle of 20 degrees while averaging over the azimuth angle. The astrophysical and observational parameters used in simulating the datasets, as well as the chosen IRFs and spectral models, are summarised in Tables 1 and 2.
@@ 3 Derivation of ALP sensitivity with the likelihood-ratio test
We consider an array of $10\times 10$ logarithmically spaced points over the ALP parameter space, with $m_{a}\in[0.1,1000]\,{\rm neV}$ and $g_{a\gamma}\in[0.03,7]\times 10^{-11}\,{\rm GeV}^{-1}$. The choice of the mass interval is commonly adopted in similar studies, and roughly corresponds to the region where the critical energy for the strong mixing regime falls within the CTAO energy range Meyer et al. (2014). Such energy value is defined as that above which the ALP-photon oscillation probability in a constant magnetic field becomes maximal and almost independent of energy Meyer et al. (2014); Mirizzi et al. (2008); Bassan et al. (2010); De Angelis et al. (2011); Galanti and Roncadelli (2022). On the other hand, the upper bound on $g_{a\gamma}$ is motivated by the most recent experimental limit set by the CAST helioscope Altenmüller et al. (2024).
Given the intrinsic spectrum of each source, we compute the photon survival probability $P_{\gamma\gamma}$ for each $(m_{a},g_{a\gamma})$ pair with gammaALPs as described in Section 2.2, and obtain the expected observed spectrum as in eq. (8). A likelihood function
$\mathcal{L}(m_{a},g_{a\gamma},\mu,b,\theta|D)=\prod_{i}\mathcal{L}(m_{a},g_{a\gamma},\mu_{i},b_{i},\theta|D_{i})$ (9)
@@
$\begin{aligned} \mathcal{L}(m_{a},&g_{a\gamma},\mu_{i},b_{i},\theta|D)=\\ &\frac{(\mu_{i}+\alpha\,b_{i})^{n_{{\rm ON},i}}}{n_{{\rm ON},i}!}e^{-(\mu_{i}+\alpha\,b_{i})}\frac{b_{i}^{n_{{\rm OFF},i}}}{n_{{\rm OFF},i}!}e^{-b_{i}}\end{aligned}$ (10)
@@
Following ref. Abe et al. (2024a), we then define the test statistic
${\rm TS}(m_{a},g_{a\gamma}|D)=-2\ln\frac{\mathcal{L}(m_{a},g_{a\gamma},\hat{\mu},\hat{b},\hat{\theta}|D)}{\hat{\mathcal{L}}(D)}\,$, (11)
@@
[…]
which corresponds to $2\ln\mathcal{L}$ after neglecting terms that do not depend on $\mu_{i}$ or $b_{i}$. This quantity is computed and returned by Gammapy when fitting each spectral model to the dataset $D$. Therefore, eq. (11) is equivalent to
$\begin{aligned} {\rm TS}(m_{a},&g_{a\gamma}|D)\\ &={\rm WSTAT}(m_{a},g_{a\gamma}|D)-{\rm WSTAT}_{\rm min}(D)\,.\end{aligned}$ (13)
@@
Figure 3: Example likelihood-ratio TS distribution obtained for the baseline state of Mrk 501
For each source (considered in both baseline and flaring states) we simulate 100 observations without including ALP effects, following the prescriptions outlined in Section 2.3, and compute the average TS for each $(m_{a},g_{a\gamma})$ pair. In order to obtain the statistical significance for excluding the ALP hypothesis with parameters $(m_{a},g_{a\gamma})$, one then needs to compare the obtained average, denoted with ${\rm TS_{0}}(m_{a},g_{a\gamma})$, against the TS distribution expected under the ALP hypothesis. As has been argued several times, Wilks’ theorem Wilks (1938) cannot be applied to the case of ALPs Meyer and Conrad (2014); Abe et al. (2024a); Abdalla et al. (2021), and thus the shape of the TS distribution cannot be assumed a priori. Therefore, for each $(m_{a},g_{a\gamma})$ point and each considered source, we simulate 100 datasets including ALP effects and compute the corresponding TS values according to eq. (13). An example of a TS distribution obtained in this way is shown in Figure 3. Following ref. Abe et al. (2024a), we fit this distribution with a Gamma probability distribution function (PDF), given by
$f(x,\alpha,\beta)=\frac{\beta^{\alpha}x^{\alpha-1}e^{-\beta x}}{\Gamma(\alpha)}\,$, (14)
@@
All the statistical computations discussed above were performed using the Scipy Python package Virtanen et al. (2020).
@@ 4.1 Machine learning classifiers
A basic example of a ML classifier is the binary decision tree (BDT). A BDT can act on a dataset [^8] In machine learning literature the term “sample” would be used, whereas “dataset” would denote a set of many samples., identified by a vector of features $\mathbf{X}=(X_{1},\dots,X_{n})$, and distinguish whether it belongs to one of two classes by operating sequential binary splits on the features’ values. The splitting thresholds at each step (node) are referred to as the parameters of the tree; on the other hand, quantities such as the tree depth, maximum number of splits, etc. are called hyperparameters. During the “supervised training” of the BDT, the splits are chosen in order to minimize some loss function, by operating on datasets whose class is known a priori. Once the tree has been trained and its parameters have been fixed, it can be used to classify previously “unseen” test datasets. There are several methods to evaluate the performance of a ML classifier during the training phase, e.g. evaluating its accuracy and related metrics with train-test splitting and cross-validation. This ensures that a similar performance will be achieved on new data as well. A vast literature of textbooks and papers exists on the many scientific applications of ML classifiers and BDTs; see e.g. ref. Raschka and Mirjalili (2017) for a basic introduction.
Random forests (RFs) are an example of ensemble classifiers, since, as the name suggests, they are based on averaging a large number of binary decision trees. In order to avoid overfitting, i.e. classifiers performing “too well” on training data but badly on test data, it is necessary that the single BDTs stop after few splits (it is said that they are “weak” classifiers). BDTs, RFs and several other ML classification algorithms are implemented in the open-source Python package Scikit-learn Pedregosa et al. (2011). When one of these algorithms is trained to distinguish between two classes and then applied to some sample, it typically returns a set of scores indicating how likely the given sample is to belong to each of the classes.
Machine learning classifiers such as RFs are widely used in Cherenkov telescope data analysis, specifically in steps that involve the so-called gamma-hadron separation Bock et al. (2004); Albert et al. (2008); Ohm et al. (2009). Here, the atmospheric shower images are parametrized by a set of quantities (the Hillas parameters Hillas (1985)) which are then used as features to distinguish between showers initiated by hadrons and by gamma-ray photons.
XGBoost (eXtreme Gradient Boosting) Chen and Guestrin (2016) is a similar algorithm to the RF, which leverages a technique called gradient boosting to efficiently minimise the loss function. Implementations of the XGBoost algorithm are available in several programming languages, including Python [^9] https://xgboost.readthedocs.io/en/stable/, and generally exhibit reasonably good performance “out of the box”, with minimal tweaking compared to RFs.